OpenStudy (anonymous):

If a plane can travel 510 mph with the wind and 430 mph against the wind, find the speed of the plane without a wind and the speed of the wind.

OpenStudy (amistre64):

510 = a + w 430 = a - w

OpenStudy (amistre64):

so you airplane is the average of the 2 ....

OpenStudy (amistre64):

do you know how to find the middle value between 2 numbers?

OpenStudy (amistre64):

thats one way of seeing it yes; the winds cancel

OpenStudy (amistre64):

thats the way we prefer it yes ....

OpenStudy (amistre64):

a+a = 2a

OpenStudy (amistre64):

510 = a + w 430 = a - w ----------- 940 = 2a

OpenStudy (amistre64):

thats correct :)

OpenStudy (amistre64):

thats what i see it as

OpenStudy (amistre64):

youre welcome